OpenStudy (anonymous):

How do I change: y + 3 = 3/4 (x + 1) into Y=mx + b form to graph it?

OpenStudy (anonymous):

why dont you try to multiply the 3/4 in (x+1) first

OpenStudy (anonymous):

y+3=3/4x+3/4 y=3/4x+3/4-3

OpenStudy (whpalmer4):

All you have to do is solve the equation for \(y\) and simplify the right hand side...

OpenStudy (anonymous):

try to do the rest by yourself.